Enrodes is a monopoly provider of residential electricity in a region of northern Michigan. Total demand by its 2 million households is Qd = 1,400 - 1P, and Enrodes can produce electricity at a constant marginal cost of $4 per megawatt hour. Consumers in this region of Michigan have recently complained that Enrodes is charging too much for its services. In fact, a few consumers are so upset that they’re trying to form a coalition to lobby the local government to regulate the price Enrodes charges.

If all the consumers of this region joined the coalition against Enrodes, how much would each consumer be willing to spend to lobby the local government to regulate Enrodes’s price?

Instruction: Enter your response rounded to the nearest penny (two decimal places).

$  
Answer #1

There is a monopoly provider of residential electricity where demand is Qd = 1,400 - 1P

This implies that inverse demand is P = 1400 - Q and marginal revenue is MR = 1400 - 2Q

The monopoly can produce electricity at a constant marginal cost = $4 per megawatt hour.

Current market outcome has MR = MC which gives 1400 - 2Q = 4 or Q= 698

This gives P = $702 per megawatt hour.

At this stage consumer surplus is 0.5*(max price - current price)*qty

= 0.5*(1400 - 702)*698 = $243,602

Consumers coalition to lobby the local government to regulate the price monopoly charges.

If regulated monopoly will charge a price of $4.

At this price, quantity is 1400 - 4 = 1396 units

Consumer surplus at this price is 0.5*(1400 - 4)*1396 = $974,408

Hence consumer can lobby for an amount which is the difference between two consumer surpluses

= $974,408 - $243,602

= $730,806

This is the required answer.